This morning I've been having problems in trying to lock the X arm.
That doesn't seem to have solved the problem. The X arm can get locked but TRX slowly moves between 0.2 and 1.
The X arm is now locked with TRX stable at ~1.
I think earlier on today I was having problems with running the alignment scripts from op540. Now I'm controlling the IFO from Rosalba and I can easily and stably lock all degrees of freedom.
I needed the X arm to be locked to align the auxiliary beam of the AbsL experiment to the IFO. To further stabilize TRX I increased the loop gain from 1 to 1.5.
Now the auxiliary beam is well aligned to the IFO and the beat is going through the PRC. I'm finally ready to scan the recycling cavity.
I also changed the gain of the PRC loop from -0.1 to -0.5.